OpenStudy (anonymous):

–6p + 7 = 3(2p – 3) – 4(–10 + 4p)

OpenStudy (anonymous):

assuming to find p -6p + 7 = 6p-9 +40-16p 12p-16p +31 = 7 -4p = -24 p = 6